# The numerical Amitsur group

### v1

@@ 1. Introduction
Let $J$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms. In this paper, the natural examples of such $J$ are subgroups of the automorphism group $\operatorname{Aut}(X)$, but many of our results apply more generally. Suppose that the Euler-Poincaré characteristic is $J$-invariant as a function of $\operatorname{Pic}(X)$; in other words $\chi(gD)=\chi(D)$ for every line bundle $D$ and element $g\in J$.
@@
The rest of the paper is structured as follows. In Section 2, we establish basic results about linearization and the Amitsur subgroup. In Section 3, we consider numerical polynomials and prove several technical results that will be used in subsequent sections. In Section 4, we explore how the Euler-Poincaré characteristic can be used to bound exponents of elements in the Amitsur group; in particular, we prove Theorem 1.3. In Section 5, we introduce the numerical Amitsur group and prove Theorems 1.1 and 1.2. In Section 6, we determine the Amitsur subgroups of toric varieties and prove Theorem 1.4. In Section 7, we work out several toric examples in detail; this serves both as a demonstration of the theory and to illustrate some of its subtleties.
@@ 2.1. Linearizations
[…]
@@
A locally free sheaf $\mathcal{E}$ is $G$-invariant if $g^{\ast}\mathcal{E}\cong\mathcal{E}$ for every $g\in G$. In the case where $\mathcal{L}$ is an invertible sheaf, $\mathcal{L}$ is $G$-invariant if and only if $[\mathcal{L}]\in\operatorname{Pic}(X)^{G}=\operatorname{Pic}(X)^{\operatorname{AutP}(X,G)}$.
@@
Suppose $\mathcal{L}$ is a $G$-invariant invertible sheaf on $X$. The Amitsur period of $\mathcal{L}$, denoted $\mathsf{m}(X,G;\mathcal{L})$, is the exponent of the element $\partial(\mathcal{L})$ in the group $H^{2}(G,\mathbb{C}^{\times})$. We use the shorthand $\mathsf{m}(\mathcal{L})$ when the $G$-variety is clear. The Amitsur period of the $G$-variety $X$, denoted $\mathsf{m}(X,G)$, is the least common multiple of $\mathsf{m}(X,G;\mathcal{L})$ over all line bundles $[\mathcal{L}]\in\operatorname{Pic}(X)^{G}$.
@@
If $[\mathcal{L}]$ is in $\operatorname{Pic}(X)^{G}$ and there exists a $G_{\mathcal{L}}$-subrepresentation $V$ of $H^{i}(X,\mathcal{L})$ of dimension $d$, then $d\partial(\mathcal{L})=0$. In other words, $\mathsf{m}(\mathcal{L})$ divides $d$.
@@
We have a $d$-dimensional representation $V$ of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ corresponds to multiplication by scalar matrices. The representation $\Lambda^{d}V$ is a $1$-dimensional representation of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ is the power map $x\mapsto x^{d}$. Therefore, we have a homomorphism $G_{\mathcal{L}}\to\mathbb{C}^{\times}$ which factors through $G_{d\mathcal{L}}$ in the diagram above. The resulting quotient morphism $G_{d\mathcal{L}}\to\mathbb{C}^{\times}$ gives the desired splitting. ∎
@@ 2.2. Amitsur subgroup
If $\varphi:H\to G$ is a group homomorphism then the kernel of the homomorphism $\operatorname{Am}(X,G)\to\operatorname{Am}(X,H)$ is the set of classes $[\mathcal{L}]$ where $\mathcal{L}$ is a $G$-invariant line bundle on $X$ such that $\varphi$ factors through $G_{\mathcal{L}}\to G$.
@@ 3.1. Univariate Polynomials
It follows immediately from the definitions that $\gcd\{f(k)\}$ divides $\gcd\{kf(k)\}$. Now we apply Corollary 3.7. Since the equations are homogeneous, we may safely divide by $\gcd\{f(k)\}$ and assume that $a_{0},\ldots,a_{n}$ have no common factor. Let $p$ be a prime and $s$ be a positive integer such that $p^{s}$ divides $\gcd\{kf(k)\}$. It remains to prove that $p^{s}\leq n+1$.
Suppose otherwise; that $p^{s}>n+1$. Consider $10<k\leq n+$ 1. Write $k=p^{r}m$ where $r$ is a non-negative integer and $m$ is a positive integer coprime to $p$. Since $k<p^{s}$, we have $r<s$. Therefore, we have the following:
@@ 4. Upper bounds on Amitsur periods
Suppose $X$ is a smooth projective variety with an action of a finite group $G$. Recall that, if $\mathcal{L}$ is $G$-invariant line bundle, then the Amitsur period $\mathsf{m}(X,G;\mathcal{L})=\mathsf{m}(\mathcal{L})$ is the exponent of $\partial[\mathcal{L}]$ in the Amitsur group $\operatorname{Am}(X,G)$. The value $\mathsf{m}(X,G)$ is the exponent of the entire group $\operatorname{Am}(X,G)$.
@@
By Proposition 2.7, we have $\chi(\mathcal{L})\partial(\mathcal{L})=0$ for every $G$-invariant line bundle $\mathcal{L}$. Using additive notation with a divisor $D$ representing the class of $\mathcal{L}$, this means that
$\left\langle\chi(kD)k[D]\ \middle|\ k\in\mathbb{Z}\right\rangle$
is a subgroup of the image $\overline{\operatorname{Pic}}(X,G)$ of $\operatorname{Pic}(X,G)$ in $\operatorname{Pic}(X)^{G}$. This subgroup is generated by $\gcd\{kf(k)\}[D]$. ∎
@@ 4.1. A uniform bound
The choice of finite group is irrelevant. We consider the polynomial $f(k)=\chi(kD)$ where $D$ represents a $G$-invariant divisor class. Observe that $f(0)=\chi(\mathcal{O}_{X})=(1+(-1)^{n}p_{a})$. Therefore $\gcd\{f(k)\}$ divides $f(0)$. The result follows immediately from Lemmas 3.8 and 4.1. ∎
@@
For any positive integer $m$, there exists a choice of finite group $G$ such that $\operatorname{Am}(\mathbb{P}^{m-1},G)\cong\mathbb{Z}/m\mathbb{Z}$. If $n\geq m-1$, then we may construct $X\cong\mathbb{P}^{m-1}\times\mathbb{P}^{n-m+1}$ where $G$ acts trivially on $\mathbb{P}^{n-m+1}$. We see that $X$ has dimension $n$ and $\operatorname{Am}(X,G)\cong\mathbb{Z}/m\mathbb{Z}$ has exponent $m$. This construction works for any $m\leq n+1$; in particular for any prime power $\leq n+1$. Therefore, for varieties of genus 00, the bound from Theorem 4.2 is multiplicatively sharp.
@@ 4.2. Bounds on the Amitsur period for specific divisors
Suppose $X$ is a curve with genus $p_{a}$ and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@
Suppose $X$ is a surface with arithmetic genus 00 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@
Suppose $X$ is a threefold with arithmetic genus 00 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@ 5. Numerical Amitsur group
Suppose $X$ is a smooth projective variety. Let $G$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms such that $\chi(D)=\chi(g^{\ast}D)$ for all $g\in G$ and $[D]\in\operatorname{Pic}(X)$. The reader should imagine that $G$ is $\operatorname{AutP}(X,G)$ from above, although we do not necessarily assume that $G$ comes from an action on $X$.
We use the following shorthand for orbit sums. For $[D]\in\operatorname{Pic}(X)$, write
@@
(5.1) $0\to\operatorname{Pic}^{\chi}(X,G)\to\operatorname{Pic}(X)^{G}\to\operatorname{Am}^{\chi}(X,G)\to 0$
@@
A divisor class $[D]$ is numerically $G$-split if and only if $[D]\in\operatorname{Pic}^{\chi}(X,G)$. (We anticipate that $\operatorname{Pic}^{\chi}(X,G)$ will also be useful in the arithmetic setting, which explains the terminology “split” instead of “linearizable.”)
@@
$[D]=\chi(E)\sum_{[F]\in G\cdot[E]}[F]$
for some divisor class $[E]\in\operatorname{Pic}(X)$. Let $H$ be the stabilizer of $[E]$ in $\operatorname{Pic}(X)$. Then $\chi(E)[E]$ is $H$-linearizable by Proposition 2.6. Thus $[D]$ is $G$-linearizable by Lemma 5.3 below. ∎
@@
Suppose $H$ is the stabilizer in $G$ of $[D]\in\operatorname{Pic(X)}$. If $D$ is $H$-linearizable, then $\Sigma_{G}(D)$ is $G$-linearizable.
@@
Observe that the determinant of $\mathcal{F}$ has the same class as $[D]$. If $\mathcal{F}$ has a linearization, then so must its determinant. Observe that the determinant is $\Sigma_{G}(D)$. Thus, it remains to exhibit a $G$-linearization on $\mathcal{F}$. We do so using a construction inspired from the “induced representation” from ordinary representation theory.
We change to the language of total spaces. Let $L$ (resp. $F$) be the total space of $\mathcal{L}$ (resp. $\mathcal{F}$) and let $\pi:L\to X$ be the projection. Let $L^{s}$ be the $s$-fold product of $L$ (as a variety) and let $p_{i}:L^{s}\to L$ denote the $i$th projection. For each $g\in G$ and index $i$ there is a unique index $g(i)$ and element $h_{g,i}\in H$ such that $gt_{i}=t_{g(i)}h_{g,i}$. We define an action of $G$ on $L^{s}$ by requiring that
@@
for each index $i$. One checks that this gives a well-defined $G$-action.
@@
The group $\operatorname{Pic}(X)^{H}$ is a finitely generated free abelian group and $\chi$ is a polynomial in the generators. Thus, by Proposition 3.3 there exists a finite set $S_{H}$ such that
@@
If $[D]\in\operatorname{Pic}(X)$ and $H$ is a subgroup of $K=\operatorname{Stab}_{G}([D])$, then
@@ 6. Toric varieties
(6.1) $0\to M\to\operatorname{TDiv}(X)\to\operatorname{Pic}(X)\to 0$
where $\operatorname{TDiv}(X)\cong\mathbb{Z}^{\Sigma(1)}\cong\mathbb{Z}^{r}$ is a free abelian group with basis indexed by the rays $\rho_{1},\ldots,\rho_{r}$ of $\Sigma$. Each ray corresponds to an irreducible $T$-invariant divisor of $X$. Let $D_{1},\ldots,D_{s}\in\operatorname{TDiv}(X)$ be a system of distinct representatives for the divisor classes in $\operatorname{Pic}(X)$ corresponding to these rays. For each $1\leq i\leq s$, let $n_{i}$ denote the number of rays in the class $[D_{i}]$.
From [Cox95], we recall the Cox ring and its connection to the automorphism group of $X$. The Cox ring $\operatorname{Cox}(X)$ of $X$ is a $\operatorname{Pic}(X)$-graded polynomial ring $\mathbb{C}[x_{1},\ldots,x_{r}]$ where each $x_{i}$ has the grading of the corresponding class in $\operatorname{Pic}(X)$. For every divisor class $[D]\in\operatorname{Pic}(X)$, we have
@@
where $U$ is a unipotent group. Each $\operatorname{GL}_{n_{i}}(\mathbb{C})$ acts on the linear span $V_{i}$ of the $n_{i}$ variables $\{x_{j}\}$ with degree $[D_{i}]$. The group $\operatorname{AutP}(X)$ acts by permuting the subspaces $\{V_{i}\}$. We have the description
@@
[…]
and define
@@
$0\to M^{W}\to\operatorname{TDiv}(X)^{W}\to\operatorname{Pic}(X)^{J}\to H^{1}(W,M)\to 0$
@@
Let $H=\prod_{i=1}^{s}S_{n_{i}}$ denote the kernel of $W\to J$. The group $H$ acts on the set of rays $\{\rho_{i}\}$ by all permutations that leave invariant the divisor classes. Therefore, the $H$-orbit of $D_{i}$ is equal to the set of basis elements of $\operatorname{TDiv}(X)$ with degree $[D_{i}]$. Consequently,
@@
The divisor classes are all distinct, so the stabilizer of $J$ acting on $\beta_{i}$ is the same as that of $[D_{i}]$ in $\operatorname{Pic}(X)$. Therefore
$[\Sigma_{W}(D_{i})]=[\Sigma_{J}(\beta_{i})]=\Sigma_{J}([\beta_{i}])=n_{i}\Sigma_{J}([D_{i}]).$
@@
Observe that $R$ acts on $\operatorname{Pic(X)}$ through $\operatorname{Aut}(X)$. Let $\widetilde{H}$ be the stabilizer in $\widetilde{G}$ of $[D_{i}]$ and let $H:=\widetilde{H}/S$ be its image in $\operatorname{Aut}(X)$. Since $R$ acts on the Cox ring, we see that $\widetilde{H}$ factors through the lifting group $H_{[D_{i}]}$ by Lemma 6.1. Observe that $\widetilde{H}$ leaves invariant the subspace $V_{i}$ of $\operatorname{Cox}(X)_{[D_{i}]}$ spanned by the variables $x_{j}$ with degree $[D_{i}]$. Therefore, there is a $n_{i}$-dimensional subrepresentation of the lifting group $H_{[D_{i}]}$ in $H^{0}(X,\mathcal{O}_{X}(D_{i}))$. Therefore, $n_{i}[D_{i}]$ is $H$-linearizable by Proposition 2.6.
@@
as subgroups of $\operatorname{Pic}(X)^{J}$. Let $\mathcal{L}$ be a $G$-linearizable line bundle. We will prove that $\mathcal{L}$ is linearly equivalent to a divisor from $\operatorname{TDiv}(X)^{W}$, which will establish that $[\mathcal{L}]$ is in $\operatorname{Pic}^{T}(X,J)$ via Lemma 6.3.
@@
where $D_{\rho_{1}},\ldots,D_{\rho_{r}}$ are the divisors corresponding to the rays of $X$. Since $-K_{X}$ is ample, using a vanishing theorem of M. Mustaţă [CLS11, Theorem 9.3.7] we see that

### v2

@@ 1. Introduction
Let $J$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms. Suppose that the Euler-Poincaré characteristic is $J$-invariant as a function of $\operatorname{Pic}(X)$; in other words $\chi(gD)=\chi(D)$ for every line bundle $D$ and element $g\in J$. This is not a very restrictive assumption since it holds for all actions coming from automorphisms of $X$ and all Galois group actions coming from the arithmetic setting. However, it excludes totally arbitrary actions such as $\mathcal{O}(m)\mapsto\mathcal{O}(-m)$ on $\operatorname{Pic}(\mathbb{P}^{n})$ that do not naturally arise from geometry.
@@
The rest of the paper is structured as follows. In Section 2, we establish basic results about linearization and the Amitsur subgroup. In Section 3, we consider numerical polynomials and prove several technical results that will be used in subsequent sections. In Section 4, we explore how the Euler-Poincaré characteristic can be used to bound orders of elements in the Amitsur group; in particular, we prove Theorem 1.3. In Section 5, we introduce the numerical Amitsur group and prove Theorems 1.1 and 1.2. In Section 6, we determine the Amitsur subgroups of toric varieties and prove Theorem 1.4. In Section 7, we work out several toric examples in detail; this serves both as a demonstration of the theory and to illustrate some of its subtleties.
@@ 2.1. Linearizations
$\operatorname{AutP}(X,G):=\operatorname{im}\left(G\to\operatorname{Aut(X)}\xrightarrow{\alpha}\operatorname{Aut}(\operatorname{Pic}(X))\right).$
@@
A locally free sheaf $\mathcal{E}$ is $G$-invariant if $g^{\ast}\mathcal{E}\cong\mathcal{E}$ for every $g\in G$. In the case where $\mathcal{L}$ is an invertible sheaf, $\mathcal{L}$ is $G$-invariant if and only if $\mathcal{L}]\in\operatorname{Pic}(X)^{G}=\operatorname{Pic}(X)^{\operatorname{AutP}(X,G)}$.
@@
Suppose $\mathcal{L}$ is a $G$-invariant invertible sheaf on $X$. The Amitsur period of $\mathcal{L}$, denoted $\mathsf{m}(X,G;\mathcal{L})$, is the order of the element $\partial(\mathcal{L})$ in the group $H^{2}(G,\mathbb{C}^{\times})$. We use the shorthand $\mathsf{m}(\mathcal{L})$ when the $G$-variety is clear. The Amitsur period of the $G$-variety $X$, denoted $\mathsf{m}(X,G)$, is the least common multiple of $\mathsf{m}(X,G;\mathcal{L})$ over all line bundles $\mathcal{L}]\in\operatorname{Pic}(X)^{G}$.
@@
If $\mathcal{L}$] is in $\operatorname{Pic}(X)^{G}$ and there exists a $G_{\mathcal{L}}$-subrepresentation $V$ of $H^{i}(X,\mathcal{L})$ of dimension $d$, then $d\partial(\mathcal{L})=0$. In other words, $\mathsf{m}(\mathcal{L})$ divides $d$.
@@
We have a $d$-dimensional representation $V$ of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ corresponds to multiplication by scalar matrices. The representation $\Lambda^{d}V$ is a $1$-dimensional representation of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ is the power map $x\mapsto x^{d}$. Therefore, we have a homomorphism $G_{\mathcal{L}}\to\mathbb{C}^{\times}$ which factors through $G_{d\mathcal{L}}$ in the diagram above. The resulting quotient morphism $G_{d\mathcal{L}}\to\mathbb{C}^{\times}$ is a retract for the morphism $\alpha:\mathbb{C}^{\times}\to G_{d\mathcal{L}}$. The kernel of this retract gives the desired splitting $G\to G_{d\mathcal{L}}$. ∎
@@ 2.2. Amitsur subgroup
If $\varphi:H\to G$ is a group homomorphism then the kernel of the homomorphism $\operatorname{Am}(X,G)\to\operatorname{Am}(X,H)$ is the set of classes $\mathcal{L}$] where $\mathcal{L}$ is a $G$-invariant line bundle on $X$ such that $\varphi$ factors through $G_{\mathcal{L}}\to G$.
@@ 3.1. Univariate Polynomials
It follows immediately from the definitions that $\gcd\{f(k)\}$ divides $\gcd\{kf(k)\}$. By Lemma 3.5, we see that
$\gcd\left\{\frac{a_{0}}{\gcd\{f(k)\}},\cdots\frac{a_{n}}{\gcd\{f(k)\}}\right\}=1;$
thus, we may assume $a_{0},\ldots,a_{n}$ have no common factor. Let $p$ be a prime and $s$ be a positive integer such that $p^{s}$ divides $\gcd\{kf(k)\}$. It remains to prove that $p^{s}\leq n+1$.
Suppose otherwise; that $p^{s}>n+1$. Consider $0<k\leq n+1$. Write $k=p^{r}m$ where $r$ is a non-negative integer and $m$ is a positive integer coprime to $p$. Since $k<p^{s}$, we have $r<s$. Applying Corollary 3.7, we have the following:
@@ 4. Upper bounds on Amitsur periods
Suppose $X$ is a smooth projective variety with an action of a finite group $G$. Recall that, if $\mathcal{L}$ is $G$-invariant line bundle, then the Amitsur period $\mathsf{m}(X,G;\mathcal{L})=\mathsf{m}(\mathcal{L})$ is the order of $\partial[\mathcal{L}]$ in the Amitsur group $\operatorname{Am}(X,G)$. The value $\mathsf{m}(X,G)$ is the exponent of the entire group $\operatorname{Am}(X,G)$.
@@
Let us use additive notation with a divisor $D$ representing the class of $\mathcal{L}$; thus $f(k)=\chi(kD)$ for every integer $k\in\mathbb{Z}$. By Proposition 2.7, we have $\chi(\mathcal{L})\partial(\mathcal{L})=0$ for every $G$-invariant line bundle $\mathcal{L}$. Thus
$\partial(f(k)kD)=\chi(kD)\partial(kD)=0$
for every integer $k$. In other words, $f(k)k[D]$ is $G$-linearizable for every $k\in\mathbb{Z}$. This means that
$A=\left\langle kf(k)[D]\ \middle|\ k\in\mathbb{Z}\right\rangle$
is a subgroup of the image $\overline{\operatorname{Pic}}(X,G)$ of $\operatorname{Pic}(X,G)$ in $\operatorname{Pic}(X)^{G}$. Now $A$ is a subgroup of $\langle[D]\rangle\cong\mathbb{Z}$. Thus $A$ is generated by $\gcd\{kf(k)\}[D]$. Since
$\overline{\operatorname{Pic}}(X,G)\cap\langle[D]\rangle=\left\langle\mathsf{m}(D)[D]\right\rangle,$
we conclude that $\mathsf{m}(D)$ divides $\gcd\{kf(k)\}$. ∎
@@ 4.1. A uniform bound
The choice of finite group is irrelevant. We consider the polynomial $f(k)=\chi(kD)$ where $D$ represents a $G$-invariant divisor class. Observe that $f(0)=\chi(\mathcal{O}_{X})=(1+(-1)^{n}p_{a})$. Therefore $\gcd\{f(k)\}$ divides $(1+(-1)^{n}p_{a})$. From Lemma 3.8, $\gcd\{kf(k)\}$ divides $\gcd\{f(k)\}\operatorname{lcm}\{1,\dots,n+1\}$, which, in turn, divides $(1+(-1)^{n}p_{a})\operatorname{lcm}\{1,\dots,n+1\}$. Now from Lemma 4.1, $\mathsf{m}(X,G)$ divides $\gcd\{kf(k)\}$. Thus, $\mathsf{m}(X,G)$ divides $(1+(-1)^{n}p_{a})\operatorname{lcm}\{1,\dots,n+1\}$. ∎
@@
For any positive integer $m$, there exists a choice of finite group $G$ such that $\operatorname{Am}(\mathbb{P}^{m-1},G)\cong\mathbb{Z}/m\mathbb{Z}$. If $n\geq m-1$, then we may construct $X\cong\mathbb{P}^{m-1}\times\mathbb{P}^{n-m+1}$ where $G$ acts trivially on $\mathbb{P}^{n-m+1}$. We see that $X$ has dimension $n$ and $\operatorname{Am}(X,G)\cong\mathbb{Z}/m\mathbb{Z}$ has exponent $m$. This construction works for any $m\leq n+1$; in particular for any prime power $\leq n+1$. Therefore, for varieties of genus 0, the bound from Theorem 4.2 is multiplicatively sharp.
@@ 4.2. Bounds on the Amitsur period for specific divisors
Suppose $X$ is a smooth curve with genus $p_{a}$ and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@
Suppose $X$ is a smooth surface with arithmetic genus 0 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@
Suppose $X$ is a smooth threefold with arithmetic genus 0 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
@@ 5. Numerical Amitsur group
Suppose $X$ is a smooth projective variety. Let $G$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms such that $\chi(D)=\chi(g^{\ast}D)$ for all $g\in G$ and $D]\in\operatorname{Pic}(X)$. The reader should imagine that $G$ is $\operatorname{AutP}(X,G)$ from above, although we do not necessarily assume that $G$ comes from an action on $X$.
We use the following shorthand for orbit sums. For $D]\in\operatorname{Pic}(X)$, write
@@
(5.1) $0\to\operatorname{Pic}^{\chi}(X,G)\to\operatorname{Pic}(X)^{G}\to\operatorname{Am}^{\chi}(X,G)\to 0.$
@@
A divisor class $[D]$ is numerically $G$-split if and only if $D]\in\operatorname{Pic}^{\chi}(X,G)$. (We anticipate that $\operatorname{Pic}^{\chi}(X,G)$ will also be useful in the arithmetic setting, which explains the terminology “split” instead of “linearizable.”)
@@
$D]=\chi(E)\sum_{[F]\in G\cdot[E]}[F]$
for some divisor class $E]\in\operatorname{Pic}(X)$. Let $H$ be the stabilizer of $[E]$ in $\operatorname{Pic}(X)$. Then $\chi(E)[E]$ is $H$-linearizable by Proposition 2.6. Thus $[D]$ is $G$-linearizable by Lemma 5.3 below. ∎
@@
Suppose $H$ is the stabilizer in $G$ of $D]\in\operatorname{Pic(X)}$. If $D$ is $H$-linearizable, then $\Sigma_{G}(D)$ is $G$-linearizable.
@@
If $\mathcal{F}$ has a linearization, then so must its determinant. Observe that the determinant is $\Sigma_{G}(D)$. Thus, it remains to exhibit a $G$-linearization on $\mathcal{F}$. We do so using a construction inspired from the “induced representation” from ordinary representation theory.
We change to the language of total spaces. Let $L$ (resp. $F$) be the total space of $\mathcal{L}$ (resp. $\mathcal{F}$) and let $\pi:L\to X$ be the projection. Since $\mathcal{L}$ is $H$-linearizable, there exists an action on $H$ on $L$ such that $\pi$ is $H$-equivariant. Let $L^{s}$ be the $s$-fold product of $L$ (as a variety) and let $p_{i}:L^{s}\to L$ denote the $i$th projection. For each $g\in G$ and index $i$ there is a unique index $g(i)$ and element $h_{g,i}\in H$ such that $gt_{i}=t_{g(i)}h_{g,i}$. We define an action of $G$ on $L^{s}$ by requiring that
@@
for each index $i$. One checks that this gives a well-defined $G$-action since the $H$-action is well-defined.
@@
The group $\operatorname{Pic}(X)^{H}$ is a finitely generated free abelian group. Thus, for each subgroup $H$, the restriction of $\chi$ to $\operatorname{Pic}(X)^{H}$ can be viewed as a polynomial of fixed degree in a finite number of variables. Thus, by Proposition 3.3 there exists a finite set $S_{H}$ such that
@@
If $D]\in\operatorname{Pic}(X)$ and $H$ is a subgroup of $K=\operatorname{Stab}_{G}([D])$, then
@@ 6. Toric varieties
(6.1) $0\to M\to\operatorname{TDiv}(X)\to\operatorname{Pic}(X)\to 0.$
where $\operatorname{TDiv}(X)\cong\mathbb{Z}^{\Sigma(1)}\cong\mathbb{Z}^{r}$ is a free abelian group with basis indexed by the rays $\rho_{1},\ldots,\rho_{r}$ of $\Sigma$. Each ray corresponds to an irreducible $T$-invariant divisor of $X$. Let $D_{1},\ldots,D_{s}\in\operatorname{TDiv}(X)$ be a system of distinct representatives for the divisor classes in $\operatorname{Pic}(X)$ corresponding to these rays. The relation between $r$ and $s$ is
$n_{1}+\dots+n_{s}=r$
where $n_{i}$ is the number of rays in each divisor class $D_{i}]$.
From [Cox95], we recall the Cox ring and its connection to the automorphism group of $X$. The Cox ring $\operatorname{Cox}(X)$ of $X$ is a $\operatorname{Pic}(X)$-graded polynomial ring $\mathbb{C}[x_{1},\ldots,x_{r}]$ where each $x_{i}$ has the grading of the corresponding class in $\operatorname{Pic}(X)$. For every divisor class $D]\in\operatorname{Pic}(X)$, we have
@@
where $U$ is a unipotent group. Each $\operatorname{GL}_{n_{i}}(\mathbb{C})$ acts on the linear span $V_{i}$ of the $n_{i}$ variables $\{x_{j}\}$ with degree $D_{i}]$. The group $\operatorname{AutP}(X)$ acts by permuting the subspaces $\{V_{i}\}$. We have the description
@@
$\operatorname{Pic}^{T}(X,J):=\langle n_{i}\Sigma_{J}([D_{i}])\mid\ 0\leq i\leq s\rangle\subseteq\operatorname{Pic}(X)^{J},$
where we recall that $n_{i}$ is the number of rays in the class $D_{i}]$, and define
@@
$0\to M^{W}\to\operatorname{TDiv}(X)^{W}\to\operatorname{Pic}(X)^{J}\to H^{1}(W,M)\to 0.$
@@
Let $H=\prod_{i=1}^{s}S_{n_{i}}$ denote the kernel of $W\to J$. The group $H$ acts on the set of rays $\{\rho_{i}\}$ by all permutations that leave invariant the divisor classes. Therefore, the $H$-orbit of $D_{i}$ is equal to the set of basis elements of $\operatorname{TDiv}(X)$ with degree $D_{i}]$. Consequently,
@@
The divisor classes are all distinct, so the stabilizer of $J$ acting on $\beta_{i}$ is the same as that of $D_{i}]$ in $\operatorname{Pic}(X)$. Therefore
$\Sigma_{W}(D_{i})]=[\Sigma_{J}(\beta_{i})]=\Sigma_{J}([\beta_{i}])=n_{i}\Sigma_{J}([D_{i}]).$
@@
Observe that $R$ acts on $\operatorname{Pic(X)}$ through $\operatorname{Aut}(X)$. Let $\widetilde{H}$ be the stabilizer in $\widetilde{G}$ of $D_{i}]$ and let $H:=\widetilde{H}/S$ be its image in $\operatorname{Aut}(X)$. Since $R$ acts on the Cox ring, we see that $\widetilde{H}$ factors through the lifting group $H_{[D_{i}]}$ by Lemma 6.1. Observe that $\widetilde{H}$ leaves invariant the subspace $V_{i}$ of $\operatorname{Cox}(X)_{[D_{i}]}$ spanned by the variables $x_{j}$ with degree $D_{i}]$. Therefore, there is a $n_{i}$-dimensional subrepresentation of the lifting group $H_{[D_{i}]}$ in $H^{0}(X,\mathcal{O}_{X}(D_{i}))$. Therefore, $n_{i}[D_{i}]$ is $H$-linearizable by Proposition 2.6.
@@
as subgroups of $\operatorname{Pic}(X)^{J}$. Let $\mathcal{L}$ be a $G$-linearizable line bundle. We will prove that $\mathcal{L}$ is linearly equivalent to a divisor from $\operatorname{TDiv}(X)^{W}$, which will establish that $\mathcal{L}$] is in $\operatorname{Pic}^{T}(X,J)$ via Lemma 6.3.
@@
where $D_{\rho_{1}},\ldots,D_{\rho_{r}}$ are the divisors corresponding to the rays of $X$. Since $K_{X}$ is ample, using a vanishing theorem of M. Mustaţă [CLS11, Theorem 9.3.7] we see that